### v1

# Learning to Disentangle Latent Reasoning Rules with Language VAEs: A Systematic Study

Abstract
Incorporating explicit reasoning rules within the latent space of language models (LMs) offers a promising pathway to enhance generalisation, interpretability, and controllability. While current Transformer-based language models have shown strong performance on Natural Language Inference (NLI) tasks, they often rely on memorisation rather than rule-based inference. This work investigates how reasoning rules can be explicitly embedded and memorised within the LMs through Language Variational Autoencoders (VAEs). We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. Performance bottleneck: In mathematical reasoning tasks using Qwen2.5 (0.5B), increasing sample count doesn’t improve performance beyond a point. Moreover, ffn layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters.
1 Introduction
Figure 1: Overview, where $(\pi,x,c)$ represents the (rule, input premise(s), conclusion). To systematically evaluate rule-based learning within a VAE framework, first, we examine three rule-based NLI tasks. Second, we formalise the hypothesis that reasoning rules can be functionally and separately encoded within the encoder’s parametric space, enabling rule learning in the latent space, grounded in the theoretical framework of neural tangent kernels. Third, we introduce an end-to-end VAE architecture, with three different latent injection setups, designed to capture coarse-grained reasoning patterns in its latent space while remaining sensitive to the lexical semantics of the input.
Encoding reasoning patterns or explicit rules within latent representations holds significant promise for enhancing the generalisation, interpretability, and controllability of neural models, which is particularly salient in the neuro-symbolic AI domain Bonnet and Macfarlane (2024); Yu et al. (2025).
Over recent years, Transformer-based language models (LMs) have achieved notable success across a variety of Natural Language Inference (NLI) tasks Yang et al. (2024); Qwen et al. (2025). Nonetheless, a growing body of research has demonstrated that these models often rely on memorisation rather than rule-based inference mechanisms, thereby constraining their capacity for generalisation and interpretability Yan et al. (2025).
Therefore, this study focuses on the question: How can reasoning rules (i.e., memorised input-output patterns [^1] Please notice that exploring the composition and generalisation over rules in the NLI task is not the target of this work. The motivation of this work is to explicit encode the memorised reasoning input-output patterns in the latent space.) be explicitly encoded within the latent space of language models?
Variational Autoencoder (VAE) Kingma and Welling (2013) provides a compelling framework for this direction. The integration of prior distribution serves as an inductive bias, enabling the model to leverage existing knowledge and providing a principled way to incorporate domain expertise Papamarkou et al. (2024). In this work, we focus on lightweight LMs (<1B parameters) as VAE decoders, allowing accessible evaluation of training dynamics, memory capacity, and model updates of the Transformer architecture Zhong et al. (2025); Morris et al. (2025). More importantly, because our primary motivation is to study latent rule encoding and separation, we fully re-train the decoder. As a result, the geometry and separation of the latent space are not constrained by the model’s size or architectures, enabling a more lightweight and controlled analysis of representation learning.
Accordingly, we propose a complete pipeline for learning NLI reasoning rules within Transformer-based language VAEs Li et al. (2020); Zhang et al. (2024c): First, we focus on three distinct rule-based reasoning tasks, each characterised by unique syntax, inference patterns and levels of granularity: Mathematical Reason Meadows et al. (2023), Syllogistic Reason Valentino et al. (2025), and Explanatory Reason Zhang et al. (2024b).
Second, following the Neural Tangent Kernel (NTK) theory Jacot et al. (2018), we claim that by supervising the inference rule information, the inference rule can be encoded in the parametric space (i.e., weight matrix) of encoder, which subsequently leads to the rule separation in the latent sentence space. This formalisation allows us to explicit encode the reasoning rule in the latent space.
Third, we propose an end-to-end VAE architecture where the latent space can encode both coarse-grained reasoning rules and fine-grained reasoning patterns that arise from semantic differences in the inputs, and evaluate three ways to inject the latent space into the LM decoder by intervening the attention network.
We conduct extensive experiments to assess the effectiveness of rule encoding and reasoning generation capabilities. The experiment illustrates the following results and findings:
Disentangled Reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space, suggesting the potential for employing rule-based NTK theory to better understand the training dynamics and internal geometry of gradient-based neural NLI model.
Prior Knowledge Injection: Different injection setups result in varying levels of reasoning performance. The optimal configuration is achieved by injecting the latent space into the query matrix of the attention network. Intuitively, injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models.
Performance Bottleneck: A case study on the mathematical reasoning task reveals a performance bottleneck in decoder-only LM, Qwen2.5 (0.5B), where increasing the number of samples per operation fails to yield performance improvements beyond a certain threshold. Additionally, we observe that ffn network is more effective than attention in maintaining the separation of reasoning rules within the learned parametric space.
To the best of our knowledge, this is the first study to explore the explicit encoding of NLI rules within the latent space of language VAEs.
2 Related Work
In this section, we review the related work around two topics: rule-based representation learning and language VAEs, to highlight the current research limitation and elucidate the motivation underlying our work.
Rule-based Representation Learning.
Encoder-only models have been employed for tasks such as mathematical operations, where the encoder learns structured transformations, as demonstrated by Valentino et al. (2024). In addition, VAE-based approaches have shown promise in tasks requiring structured reasoning, such as visual program synthesis, where the goal is to generate programs that fulfill a specified task Sun et al. (2018); Bonnet and Macfarlane (2024); van Krieken et al. (2025). Similarly, grammar-based approaches using VAEs have been applied to infer ordinary differential equation (ODE) formulas from data Yu et al. (2025). Moreover, Decoder-only models, particularly large language models (LLMs), leverage in-context learning by using demonstrations to infer and apply underlying reasoning patterns Liu et al. (2023b); Bhattamishra et al. (2024). Despite these advancements, relatively few studies have investigated rule-based learning in the context of NLI. To address this gap, we present a foundational exploration of rule-based learning applied to the NLI task.
Language VAEs.
Language VAEs have been widely applied in NLP tasks, such as style transfer tasks: modifying sentences with regard to markers of sentiment, formality, affirmation/negation Shen et al. (2020); John et al. (2019); Bao et al. (2019b); Hu and Li (2021); Vasilakes et al. (2022); Gu et al. (2022); Liu et al. (2023a); Gu et al. (2023), story generation Fang et al. (2021), dialogue generation Zhao et al. (2017), text paraphrasing Bao et al. (2019a), and textual, syntactic, semantic representation learning domain, such as syntax disentanglement Mercatali and Freitas (2021), semantic-syntax separation Zhang et al. (2024c), semantic disentanglement Silva De Carvalho et al. (2023); Zhang et al. (2024a), etc. Comparatively, we focused on the Natural Language Inference (NLI) task. To our knowledge, the application of VAEs for NLI is underexplored.
In the next section, we start by introducing the reasoning tasks and provide a formal illustration of rule-based learning through Neural Tangent Kernel theory.
3 Rule-based Learning for NLI
3.1 Natural Language Inference Rules
In this study, we investigate three distinct types of reasoning tasks: mathematical derivation Meadows et al. (2024), syllogistic reasoning Valentino et al. (2025), and explanatory reasoning Zhang et al. (2024b). By spanning a range of reasoning granularities, these tasks facilitate a systematic evaluation. A summarisation of each task and corpus is provided in Table 4.
Mathematical reasoning.
Mathematical expressions Valentino et al. (2023); Meadows et al. (2023) follow a well-defined syntactic structure and set of symbolic rules that are notoriously difficult for neural models. The dataset Meadows et al. (2023) includes seven human-annotated symbolic rules, encompassing operations such as differentiation and integration.
Syllogistic reasoning.
Syllogistic reasoning involves classical categorical logic, including four standard forms: Universal Affirmative (A) — “All A are B,” Universal Negative (E) — “No A are B,” Particular Affirmative (I) — “Some A are B,” and Particular Negative (O) — “Some A are not B.” The dataset encodes 24 valid syllogistic inference patterns, such as AA1 (Barbara).
Explanatory reasoning.
Explanatory sentences Jansen et al. (2018); Valentino et al. (2022) provide a semantically challenging yet sufficiently well-scoped scenario to evaluate the syntactic and semantic organisation of the space. Based on the EntailmentBank corpus Dalvi et al. (2021), Zhang et al. (2024b) defined ten inference types based on the transformation over Abstract Meaning Representation (AMR) graph Banarescu et al. (2013).
3.2 Rule-based Neural Tangent Kernel
Figure 2: Conceptual visualisation of rule separation in parametric space of encoder. We can quantify rule separation by computing the distance between gradient vectors associated with distinct rule categories.
This work aims to encode latent reasoning rules within a low-dimensional latent space to guide the LM decoder’s reasoning generation, leveraging a VAE framework. We frame the rule-based learning as learning the transformation from input premise(s) to output conclusion within the encoder. We formalise this framework by introducing a proposition grounded in the Neural Tangent Kernel (NTK) theory Jacot et al. (2018).
3.2.1 Latent Subspace Separation
Let $\mathcal{M}$ be a VAE model parameterised by $\theta=(\theta_{\text{enc}},\theta_{\text{dec}})$. Suppose the encoder can encode a set of symbolic inference rules $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. Then, under supervised training with rule annotations:
Proposition: the encoder’s parameters $\theta_{\text{enc}}$ induce a parametric structure in which each inference rule $\pi_{i}$ corresponds to a distinct subspace $S_{\pi_{i}}\subseteq\mathbb{R}^{D}$ of the encoder representation space.
$\forall\pi_{i},\pi_{j}\in\Pi,\,\pi_{i}\neq\pi_{j}\quad\Rightarrow\quad S_{\pi_{i}}\cap S_{\pi_{j}}\approx\emptyset$ (1)
This parametric separation directly leads to clearly delineated latent feature subspaces, each uniquely encoding a symbolic inference rule.
3.2.2 Connection to NTK Theory
NTK theory provides a rigorous theoretical framework for understanding neural network training dynamics by examining the kernel induced by gradients of the network’s outputs with respect to its parameters. It suggests that during gradient descent, the network effectively learns linear approximations within distinct functional subspaces where each approximating discrete symbolic reasoning rules in our task. Formally:
Step 1: formal setup.
Let $f_{\text{encode}}$ be the encoder function such that:
$f_{\text{encode}}:(\pi,x,c)\mapsto\mathbf{z}_{(\pi,x,c)}\in\mathcal{Z}\subseteq\mathbb{R}^{D}$ (2)
where $x$ is the input premise(s), c is the conclusion, $\pi\in\Pi$ is the reasoning rule, and $\mathcal{Z}$ is the latent representation space of dimension $D$. The training objective is to minimise the expected loss:
$\mathcal{L}(\theta_{\text{enc}})=\mathbb{E}_{(\pi,x,c)\sim\mathcal{D}}\left[-\log p_{\theta}(\pi,x,c)\right]$ (3)
Step 2: NTK as rule subspaces.
Each rule $\pi$ is explicitly embedded as part of the model input. As a result, the model effectively learns a function $f_{\theta}(\pi,x,c)$. The function $f_{\theta}$ thus jointly depends on both the content of transformation from premise(s) to conclusion and the nature of the symbolic operation to be performed.
Within the NTK framework, the similarity between two input examples of the same inference type $\pi$ is captured by the NTK as follows:
$\Theta_{\pi}(x,x^{\prime})=\nabla_{\theta}f_{\theta}(x,\pi)^{\top}\nabla_{\theta}f_{\theta}(x^{\prime},\pi)$ (4)
In here, $x$ represents $(x,c)$ pair for concision. $\nabla_{\theta}f_{\theta}(x,\pi)$ denotes the gradient of the model output with respect to its parameters, evaluated at the input $(x,\pi)$. This kernel quantifies how a parameter update from one input-output pair would affect another pair, conditioned on the shared rule.
According to NTK theory, the evolution of the model’s predictions under gradient descent training can be described by a linear kernel regression in the RKHS (Reproducing Kernel Hilbert Space) associated with $\Theta_{\pi}$. Specifically, the prediction at time $t$, $f_{t}(x,\pi)$, evolves as:
$f_{t}(x,\pi)=f_{0}(x,\pi)-\Theta_{\pi}(x,\cdot)\left[\Theta_{\pi}+\lambda I\right]^{-1}(f_{0}-c)$ (5)
where $f_{0}(x,\pi)$ is the model’s output at initialisation for each training input, $\lambda$ is a regularisation parameter, and $c$ is the vector of ground truth conclusions.
Crucially, this formulation implies that each rule $\pi$ induces a distinct kernel $\Theta_{\pi}$, which in turn defines a unique RKHS $\mathcal{H}_{\pi}$—that is, a function space within which the model’s solutions for rule $\pi$ reside. As the $\pi$ is varied, the structure of the kernel and the corresponding function space changes, reflecting the distinct reasoning behaviours associated with different inference operations. Thus, the model encodes different symbolic inference patterns in distinct, kernel-induced subspaces.
Step 3: Disjointness via kernel independence.
For two different rules, $\pi_{i}\neq\pi_{j}$, we examine the relationship between their corresponding NTKs, $\Theta_{\pi_{i}}$ and $\Theta_{\pi_{j}}$. Specifically, we are interested in the interaction between the parameter gradients induced by inputs associated with different inference types.
Consider two data points $x$ and $x^{\prime}$, possibly corresponding to different premise pairs. The NTK entry for each rule is:
$\Theta_{\pi_{i}}(x,x^{\prime})=\nabla_{\theta}f_{\theta}(x,\pi_{i})^{\top}\nabla_{\theta}f_{\theta}(x^{\prime},\pi_{i})$ (6)
and
$\Theta_{\pi_{j}}(x,x^{\prime})=\nabla_{\theta}f_{\theta}(x,\pi_{j})^{\top}\nabla_{\theta}f_{\theta}(x^{\prime},\pi_{j})$ (7)
When considering cross-rule similarities, we are interested in the inner product between the gradients for different rules:
$G_{ij}(x,x^{\prime}):=\langle\nabla_{\theta}f_{\theta}(x,\pi_{i}),\nabla_{\theta}f_{\theta}(x^{\prime},\pi_{j})\rangle$ (8)
If the rules $\pi_{i}$ and $\pi_{j}$ encode fundamentally different reasoning operations (e.g., addition vs. subtraction in Math Derivation), the gradients with respect to $\theta$ for inputs labeled with $\pi_{i}$ and those labeled with $\pi_{j}$ will tend to point in different directions in parameter space of encoder.
Under idealised training, where the data for each rule is sufficiently distinct and the network has enough capacity, the gradients for one rule will have minimal overlap with those of the other. This can be formalised by observing that:
$\langle\nabla_{\theta}f_{\theta}(x,\pi_{i}),\nabla_{\theta}f_{\theta}(x^{\prime},\pi_{j})\rangle\approx 0\qquad\text{for}~{}\pi_{i}\neq\pi_{j}$ (9)
This property implies that the parameter updates driven by examples from different rules are approximately orthogonal, meaning that training on one type will not interfere with or alter the function learned for the other type. In the language of NTK and kernel regression, this corresponds to the induced RKHS for each type, $\mathcal{H}{\pi_{i}}$ and $\mathcal{H}{\pi_{j}}$, being approximately disjoint:
$\mathcal{H}{\pi_{i}}\cap\mathcal{H}{\pi_{j}}\approx\emptyset$ (10)
In the next section, we will introduce the practical VAE architecture and optimisation function.
Base Model Math Derivations Math Derivations (OOD) Syllogistic Reasoning Explanatory Reasoning
bleu acc bleu acc bleu acc bleu
GPT2-medium 0.2019 0.0171 0.0206 0.0018 0.0108 0.0000 0.5947
Llama3-1B 0.3412 0.0257 0.0625 0.0114 0.3612 0.1200 0.3160
Qwen2-0.5B 0.4200 0.1171 0.0869 0.0128 0.8130 0.4600 0.5372
Qwen2.5-0.5B 0.6260 0.3800 0.1293 0.0185 0.9014 0.7000 0.6566
Table 1: Quantitative evaluation for decoder-only models. We can observe that Qwen2.5-0.5B demonstrates strong performance across various baselines, making it a suitable choice as the decoder component in our VAE architecture.
4 Approach
Latent space properties.
We posit that the latent space should satisfy two essential geometrical properties:
Property 1: Rule-level encoding. The latent space must capture the transformation defined by the reasoning rule $\pi$, which maps an input $x$ to a conclusion $c$, i.e., $\pi:x\rightarrow c$.
Property 2: Semantic-level encoding. The latent space should also encode the lexical semantics of the $x$, accounting for fine-grained variations in reasoning patterns that arise from semantic differences in the inputs. For example, within the Math Derivation task, under differentiation, different inputs yield distinct transformations:
[…]
By ensuring both properties, the latent space is capable of capturing both coarse-grained and fine-grained reasoning patterns.
Architecture.
We adopt a Transformer-based VAE framework, employing two distinct encoders—both instantiated with BERT Devlin et al. (2018)—to model the prior and posterior Gaussian distributions. The posterior encoder is optimised to capture the reasoning rule transformation, thereby satisfying Property 1. Simultaneously, the prior encoder serves as a regulariser, encouraging the posterior to encode fine-grained sub-rule variations grounded in the lexical semantics of the input, thereby satisfying Property 2.
Concretely, given a target $(\pi,x)$ pair, we first retrieve a set of semantic similar examples $\{(\pi,x_{1},c_{1}),\dots,(\pi,x_{N},c_{N})\}$ using BM25 retrieval (N is 12 during the experiment). These examples are input to the posterior encoder, which learns a latent representation of the reasoning transformation by averaging the latent vectors of the retrieved instances. Geometrically, averaging the latent sample vectors positions the target $x$ at the centroid of the surrounding samples within the latent space, which is naturally aligned with how information is encoded in the latent space Zhang et al. (2024a, 2025).
The decoder then utilises this aggregated representation to generate the corresponding conclusion $c$ for the target triplet $(\pi,x,c)$. In parallel, the prior encoder processes the target $(\pi,x)$ pair to predict the rule $\pi$ within the latent space via a linear classifier. By minimising the Kullback–Leibler (KL) divergence between the posterior and prior distributions, the latent space is encouraged to satisfy both geometric properties.
Latent injection.
We adopt three strategies to inject the latent variable $z$ into the decoder (e.g., Qwen2.5 Qwen et al. (2025)). Each approach modifies the attention mechanism to incorporate information from the latent space.
1. kv_add
Following Zhang et al. (2024c), the latent vector $z$ is added to both the Key ($K$) and Value ($V$) matrices in the attention network:
$\text{softmax}\left(\frac{\text{Q}[\text{z}+\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z}+\text{V}]$ (11)
2. kv_mem
Following Li et al. (2020), the latent vector $z$ is concatenated to the Key and Value matrices:
$\text{softmax}\left(\frac{\text{Q}[\text{z};\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z};\text{V}]$ (12)
3. query_add
In this setup, the latent vector $z$ is added to the Query ($Q$) matrix:
$\text{softmax}\left(\frac{[\text{Q}+\text{z}]\text{K}^{\top}}{\sqrt{d}}\right)\text{V}$ (13)
Here, $Q$, $K$, and $V$ denote the query, key, and value matrices in the attention mechanism, each with dimensions $\mathbb{R}^{\text{dim}\times\text{seq}}$, where dim is the attention dimensionality and seq is the sequence length.
Optimisation.
Finally, the model can be trained end-to-end via the evidence lower bound (ELBO) on the log-likelihood of the data $x$ Kingma and Welling (2013). To avoid the KL vanishing issue, we select the cyclical schedule to increase weights of KL $\beta$ from 0 to 1 Fu et al. (2019) and a KL thresholding scheme Li et al. (2019) that chooses the maximum between KL and threshold $\lambda$. The final objective function can be described as follows:
$\mathcal{L}_{\text{VAE}}=$ $\mathbb{E}_{q_{\phi}(z|\pi,x,c)}\Big{[}\log p_{\theta}(\pi,x,c|z)\Big{]}$ (14)
$-\beta\max\left[\lambda,\text{KL}q_{\phi}(z|\pi,x,c)||p(z|\pi,x)\right]$
[…]
Where $q_{\phi}$, $p$, and $p_{\theta}$ represent the posterior encoder, prior encoder, and decoder, respectively. cls_weight controls the strength of the classification loss. In our experiments, cls_weight is evaluated at 1.0, 0.5, and 0.1.
5 Empirical Analysis
5.1 Decoding Evaluation
Base model.
[…]
During inference, we omit $c$, allowing the model to generate the conclusion. To evaluate generation quality, we employ both the BLEU score Papineni et al. (2002) and accuracy (acc) as performance metrics. For the Math Reasoning task, in addition to the in-distribution test set, we also assess model performance on an out-of-distribution (OOD) test set, where the mathematical expressions are composed using a different set of variables.
Injection Train Setup Math Reason Math Reason (OOD) Syllogistic Reason Explanatory Reason
bleu acc bleu acc bleu acc bleu
Base model Zero-shot 0.6260 0.3800 0.1293 0.0185 0.9014 0.7000 0.6566
Few-shot 0.1596 0.0614 0.1011 0.0057 0.1534 0.0000 0.0765
Few-shot (FT) 0.1601 0.0557 0.1022 0.0028 0.1712 0.0000 0.0801
kv_add prior=False 0.7285 0.5285 0.1055 0.0200 0.4839 0.3400 0.6127
weight=1.0 0.4986 0.2128 0.0869 0.0157 0.0719 0.0000 0.2049
weight=0.5 0.4925 0.2814 0.0504 0.0185 0.4870 0.0900 0.2623
kv_mem prior=False 0.6430 0.4185 0.0985 0.0200 0.4672 0.3300 0.6313
weight=1.0 0.5255 0.2300 0.0956 0.0185 0.0000 0.0000 0.6022
weight=0.5 0.6808 0.4857 0.1130 0.0185 0.9452 0.8500 0.5987
query_add prior=False 0.6501 0.3885 0.1130 0.0200 0.9681 0.9200 0.6449
weight=1.0 0.7262 0.5057 0.1223 0.0214 0.4373 0.3300 0.6118
weight=0.5 0.8130 0.6642 0.1441 0.0185 0.9461 0.8600 0.6220
Table 2: Quantitative evaluation, where the base model is Qwen2.5-0.5B. Top two values are highlighed in blue and bold. We can observe that query_add leads to the best performance in general.
As shown in Table 1, Qwen2.5 demonstrates consistently strong performance across all tasks. Based on these results, we select Qwen2.5 as the decoder model for subsequent experiments.
VAE model.
Next, we evaluate the performance of VAE on downstream reasoning tasks, where the decoder is Qwen2.5-0.5B, the latent dimension is 32 following the same setup as Optimus Li et al. (2020). In addition, we include results from the base model trained via fine-tuning (denoted as FT) as well as inference using few-shot examples. The input format remains consistent with previous settings, with examples repeated accordingly. For few-shot selection, we employ BM25 to retrieve the most relevant samples.
As illustrated in Table 2, we can observe that injecting the latent space into the query can generally result in better performance compared with other setups. Intuitively, injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key (Finding).
Furthermore, when the base model is trained using few-shot examples, its performance declines significantly. We observe that the model repeats to generate more examples. To ensure a fair comparison across methods, we did not apply any filtering to remove these redundant outputs (Finding).
Additionally, we evaluate the reasoning performance of the VAE model trained with and without BM25-based sample selection. In the absence of BM25, training samples are randomly drawn from the corpus. As presented in Table 3, integrating BM25-based selection during training yields enhanced reasoning performance, indicating the effectiveness of relevance-guided sampling. From a geometric perspective, BM25 retrieves samples that exhibit the highest lexical similarity, effectively selecting the nearest neighbors in the latent space (Finding).
In the following section, we evaluate the model’s capability to learn and represent latent reasoning rules within the latent space.
Injection Prior BM25 bleu acc
kv_add False False 0.3691 0.0614
kv_mem False False 0.5525 0.2371
query_add False False 0.5914 0.3021
kv_add False True 0.7285 0.5285
kv_mem False True 0.6430 0.4185
query_add False True 0.6501 0.3885
Table 3: Quantitative evaluation for VAE model with or without BM25 in Math Reason task. We can observe that selecting the training sample using BM25 can help model performance.
Figure 3: Gradient heatmap for the last posterior encoder layer (query_add setup), where the left: math derivation, middle: explanatory reasoning, right: syllogistic reasoning, Top: cls_weight is 1.0, bottom: cls_weight is 0.1. We can observe that the non-diagonal values are notably close to 0 when providing higher cls_weight (the red colour elements are less scattered), suggesting that incorporating rule information during training enhances the separation of rule subspaces in the encoder’s parameter space. We provide the heatmaps of all layers in Appendix B.
5.2 Encoding Evaluation
Latent parametric space.
As we illustrated in Section 3.2 (Equation 9), by measuring the cosine similarity between gradient vectors associated with different rules, we can quantify the separability between different rule subspaces in the encoder, comparing settings with strong cls_weight = 1.0 (left) and weak cls_weight =0.1 (right) during training. As illustrated in Figure 3, when the classification weight (cls_weight) is set to 1.0, most non-diagonal values are close to zero (orthogonality). In contrast, with cls_weight set to 0.1, a greater number of non-diagonal elements exhibit higher values (the red colour elements are much more scattered). This observation suggests that explicit supervision facilitates the separation of reasoning rules within the encoder’s parameter space (Finding).
Latent sentence space.
In Section 3.2, we illustrated that rule separation in the parametric space leads to corresponding separation in the feature space. As shown in Figure 4, the sentence representations tend to form distinct clusters that reflect rule information when the classifier is given a higher cls_weight. However, when the cls_weight approaches zero, these rule-based clusters disappear. This suggests that the neural network relies more on the memorisation of lexical combination than on rule-based learning (Finding).
Figure 4: PCA visualisation for query_add injection setup, where top: cls_weight is 1.0, bottom: cls_weight is 0.1. We can observe that the model struggle to learn the rules when the weight is close to zero, indicating the neural network try to deliver reason behaviour via memorisation, rather than rule-based learning. For other injection setups, their visualisations are provided in Figure 12, 13, and 14.
Figure 5: Case study for Math Reasoning task, where left: analysing how varying the number of training samples for each operation affects the reasoning capabilities. Right: comparing the parametric rule separation between attn and ffn at the last layer in Qwen2.5-0.5B, a pretrained checkpoint with a training sample size of 4k.
5.3 Case Study for Math Reasoning
Information bottleneck.
First, we analyse how varying the number of training samples for each operation affects the reasoning capabilities of the decoder-only LM (Qwen2.5-0.5B). As illustrated in Figure 5 (left bar plot), it can be observed that when the number of samples per category exceeds 2,000, there is a noticeable decline in accuracy. This suggests that increasing the sample size may introduce greater variability or complexity, potentially disrupting the consistency of each operation. The observation also highlights a limitation of current autoregressive LMs: rather than engaging in rule-based reasoning, they tend to rely on retrieving memorised training instances embedded in their parameters Zhong et al. (2025). Explicitly injecting latent reasoning representations can help mitigate this issue (Finding).
Parametric rule separation.
Second, we assess the parametric separation of rule-based information across different model components, specifically the attention (attn) and feed-forward network (ffn) layers, using the same methodology outlined previously. We report the sum of the average diagonal values across all layers. As illustrated in Figure 5 (right bar plot), the ffn layers (right) exhibit a greater tendency to encode rule-based information compared to the attention layers, indicating a more prominent role in capturing structured reasoning patterns (Finding).
6 Conclusion and Future Work
This study serves as a foundational step in exploring the rule-based representation learning under the language VAE architecture for NLI tasks.
We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. Performance bottleneck: In mathematical reasoning tasks using Qwen2.5 (0.5B), increasing sample count doesn’t improve performance beyond a point. Moreover, ffn layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters. To investigate compositional behaviour beyond memorisation in NLI, in the future, we will focus on current diffusion-based language models, such as Mmada Yang et al. (2025) and Gemini diffusion, which offer greater flexibility for constraining and optimising the decoding process (van Krieken et al., 2025).
7 Limitations
Large scale decoder.
Since our primary motivation is to investigate latent rule separation, we focus on small-scale LM decoders. This choice allows for accessible evaluation of training dynamics, memory capacity, and model updates within the Transformer architecture. Importantly, because the decoder is fully re-trained, the resulting latent space geometry and separation are not constrained by the size or pre-training of the model, enabling a more controlled and interpretable analysis of the learned representations. However, how we can seamlessly integrate the prior rule information into the frozen LLM decoders should be explored in the future.
Composition and Generalisation.
Encoding rule information in the latent space can support compositional reasoning and enhance generalisation in Natural Language Inference (NLI) tasks. However, in this work, we primarily focus on memorisation rather than generalisation. Specifically, we define a rule as a memorised input-output pattern encoded within the latent space. While this approach facilitates controlled reasoning over known patterns, the broader generalisation capacity remains constrained by the architectural limitations of auto-regressive LM decoders. Therefore, we incorporate such prior rule information into the Query component of the attention mechanism in the LM decoder, enabling more effective retrieval through attention over the Key and Value representations. Further investigation is needed to understand how rule-based composition scales to more complex and diverse inference scenarios.
Appendix A Experimental Details
Corpus.
Table 4 describes the reasoning tasks in the experiment.
Train setup.
Max epoch: 10, lr: 3e-5, GPU device: A6000 $\times$ 2.
Task Example Num of Rules Sizes
Math Reason
Operation: differentiation
[…]
7 28k
Explanatory Reason
Operation: argument substitution
P1: a scar on the knee is a kind of scar.
P2: a scar is an acquired characteristic.
C: a scar on the knee is an acquired characteristic.
8 ~5k
Syllogistic Reason
Operation: AA1
P1: all apples are edible fruits.
P2: all edible fruits are fruits.
C: all apples are fruits.
26 ~5k
Table 4: Reasoning tasks, examples, number of rules, and dataset sizes.
Appendix B Additional Results
Parametric space visualisation.
We provide comprehensive heatmaps covering all encoder layers across the full set of tasks, including Math Reasoning: Figure 6 and 7; Explanatory Reasoning: Figure 8 and 9; Syllogistic Reasoning: Figure 10 and 11.
Figure 6: Math Reasoning: gradient heatmap, where cls_weight is 0.1.
Figure 7: Math Reasoning: gradient heatmap, where cls_weight is 1.0.
Figure 8: Explanatory Reasoning: gradient heatmap, where cls_weight is 0.1.
Figure 9: Explanatory Reasoning: gradient heatmap, where cls_weight is 1.0.
Figure 10: Syllogistic Reasoning: gradient heatmap, where cls_weight is 0.1.
Figure 11: Syllogistic Reasoning: gradient heatmap, where cls_weight is 1.0.
Feature space visualisation.
We provide the latent space visualisation across all reasoning tasks in Figure 12, 13, and 14.
Figure 12: PCA and T-SNE visualisation for Math Derivation task.
Figure 13: PCA and T-SNE visualisation for Explanatory Reasoning task.
Figure 14: PCA and T-SNE visualisation for Syllogistic Reasoning task.